OpenStudy (anonymous):

Prove using contradiction that the square root of any irrational number is irrational.

OpenStudy (perl):

Let r be irrational number Claim : sqrt( r ) is irrational: Proof : suppose sqrt(r) is rational then sqrt (r) = p/q then r = (p/q)^2 do you see the contradiction now?

OpenStudy (anonymous):

not really :(

OpenStudy (anonymous):

okay, thanks

OpenStudy (perl):

(i cleaned the proof a bit) Suppose you are given an unspecified irrational number, which we will label r. Claim : sqrt( r ) is irrational: Proof: (by contradiction): Assume the negation of the claim; i.e. assume sqrt( r ) is not irrational. This is the same thing as saying sqrt(r) is rational. Then by definition of a rational number, sqrt (r) = p/q for some integers p, q, where q is non-zero. We can square both sides of sqrt(r) = p/q. (You can justify this is a valid step using field axioms on the real numbers). sqrt(r)^2 = (p/q)^2 r = p^2/q^2 But now r is a ratio of integers, therefore r is a rational number. But we were given that r is irrational before we even did the proof. Therefore r is both rational and not rational. Contradiction . Conclusion: Since assuming the negated version of the claim led to a contradiction, the negated claim is false. That is, it is false that sqrt(r) is not irrational. Hence it is true that sqrt(r) is irrational.